Find the Reynolds number if a fluid of viscosity 0.4 Ns/m2 and relative density of 900 kg/m3
through a 20 mm pipe with a velocity of 2.5 m/s. Is the flow Laminar or Turbulent?
Note: If Reynolds < 2000 the flow is laminar
If Reynolds > 4000 the flow is turbulent
If 2000 < Re < 4000 it is transitional flow
Re =
1
Expert's answer
2020-06-08T10:30:14-0400
"Reynolds number N=\rho vl/\mu"

where "\mu=0.04Ns/m^2 , \rho=900kg/m^3, l=20mm=0.02m, v= 2.5 m/s"


"Reynolds number N=900*2.5*0.02\/0.4= 112.5"


since Reynolds number is less than 2000 so flow is laminar.